OpenStudy (anonymous):

Help to make sure I have this question done right: 5t + 12 = 4 - 3t solve for t

OpenStudy (anonymous):

t=-1

OpenStudy (anonymous):

Thats what I got, however I need to prove the solution which I don't know how to do that.

OpenStudy (anonymous):

5t-3t=2t it then becomes 3t+12=4 minus 12 from both sides and u get negative one

OpenStudy (anonymous):

Yes, I have the solution but I need to prove it.

OpenStudy (tyteen4a03):

@cahit Please don't give out answers without explanation. 5t + 12 = 4 - 3t The first step should be adding 3t to both sides (remember - opposites)

OpenStudy (anonymous):

Hi tyteen4a03, I've already gone over the answer with someone but my question now is how do I prove my solution is correct?

OpenStudy (tyteen4a03):

@joshhhunt Oh... Split the equation into two parts, the Left Hand Side (LHS) and Right Hand Side (RHS), then plug in the values and see if two sides are equal.

OpenStudy (anonymous):

I don't quite understand how I do that, could you give me an example?

OpenStudy (anonymous):

get a piece of paper and write out the problem is step one, i'll walk u through

OpenStudy (anonymous):

This is what I have for solving it: 5t + 12 = 4 - 3t 5t + 3t = 4 - 12 8t = -8 t = -1

OpenStudy (anonymous):

so there u proved it

OpenStudy (anonymous):

proving it is showing work

OpenStudy (anonymous):

I'm a little confused because it's a two part question, the first is solve for t and the second is prove it. I wouldn't of thought I put the same answer for both.

OpenStudy (tyteen4a03):

@shatikat That is not proving. Try this: 6x - 2 = 2x + 4 The correct solution is x = 3/2. Now split the two parts: LHS = 6x - 2 Sub. x = 3/2 into LHS: 6(3/2) - 2 = 7 RHS = 2x + 4 Sub. x = 3/2 into RHS: 2(3/2) + 4 = 7 = LHS Therefore the answer for x is correct.

OpenStudy (anonymous):

Ahh I think I understand, so what I do is: LHS = 5t + 12 Sub. x = -1 into LHS: 5 x -1 + 12 = 7 and then same for the RHS?

OpenStudy (tyteen4a03):

Yes. Make sure you draw a conclusion from what you got in RHS (i.e = LHS) then use this information to draw a final conclusion.